Collision Six, Senses

Art Interactive, 130 Bishop Allen Drive, Cambridge, MA Exhibit: 10/30/04 - 11/07/04, 12-6pm Opening Reception: Friday, 10/29/04, 6-9pm

The Collision Collective and Art Interactive present “Collision Six, Senses”, an experimental exploration of art and technology. Collision Six, the sixth event in the popular Collision Series, will showcase art from artists from MIT and beyond who use new technologies in their work. Eighteen pieces of art will be presented by Stefan Agamanolis, Jessica Banks, Joelle Bitton, Nell Breyer, Andrew Brooks, Ben Dalton, Jeana Frost, jackbackrack, Brian Knep, Nick Knouf, Jeff Lieberman, Dan Maynes-Aminzade, Andrew Neumann, Amanda Parkes, Nancy Paterson, James Patten, Hayes Raffle, Ramesh Raskar, Dan Roe, Fran Trainor, and William Tremblay. Collision Six will also feature Toronto based media artist Nancy Paterson’s “Stock Market Skirt”.

Collisions are a showcase of envelope-pushing artwork in an interactive workshop/laboratory format. The artwork often involves never before tried technologies, concepts and installation approaches. It is an opportunity for Collision colluders to experiment and show new ideas and techniques and to discuss their work with and gather feedback from the public. Artists will be available during the opening and weekends to speak with the public.

Curiously Strong (2004) Jessica Banks, Amanda Parkes

Cambridge, MA USA

Altoids boxes, solenoids, acrylic, copper approx. 1”(h) x 4” (w) x 50’ (l)n

the idea behind curiously strong is that of a single line of 500 dominoes, except it is implemented with altoids boxes. each box is equipped with a solenoid to open and close the lid and circuitry for activate that motion. the result is a infinitely looping sequence of closing and opening lids. the piece can be configured to the space, traveling up and over furniture and snaking around the space (e.g., onto walls, patterns on the floor, etc.). thus it is easy to integrate it with other works, taking advantage of the natural organization of the space. the

goal is to have 500 altoids boxes strung up. that is about 100 feet of distance. each box is 4 inches wide. all circuitry and connections will be within that footprint.

RAW (2003) Joelle Bitton, Stefan Agamanolis, Matthew Karau

Dublin, Ireland

Sound+still pictures+ISIS Projection-basedn

(1) Overview RAW is a system combining a tool and a process for capturing and conveying audiovisual impressions of everyday life. The tool itself incorporates a digital camera and a binaural audio recording device that captures the minute of sound before and after a picture is taken. The tool was used during a set of workshops in the African country of Mali, in August 2003 by 23 participants in Bamako, Timbuktu and Ségou. (2)

Installation The installation presents the 23 RAW records authored by these participants. The material gathered with the tool “raw” and unedited all the way from production to archive to presentation; hence the name of the project. (3)

Background The

RAW project began with the realization that, for many reasons, we don’t always have a good sense of what everyday life is like in other places in the world, and that having this sense might be helpful in improving understanding and relations between people in different cultures. Records and accounts of everyday life in our pasts and presents are often mediated by numerous third parties. We feel this mediation degrades the full sense of awareness and appreciation we could achieve of other peoples and places, above cultural stereotypes and clichés. The goal of the RAW project is to develop a new kind of recording tool, together with a method for processing and presenting the material captured with the tool, that enables a more direct, minimallymediated relationship between its user and the later audience, possibly in a far away place or time. (4)

The RAW system RAW

is a system combining a tool and a process for capturing and conveying audiovisual impressions of everyday life. The project aims to enable a relationship between the user of the tool and an audience in a different place or time with an absolute minimum of editorial mediation by a third party. The RAW tool consists of a digital still camera and a high quality digital stereo audio recorder that captures the minute of sound before and after a picture is taken. The relationship created between sound and image forms a disjoint flow and opens a new field of audiovisual expression. These previously uncaptured moments in time can be kept as personal artefacts (diaries, journals), for cultural exchanges, for story-telling production or archived for later study. Audio is recorded binaurally using high-quality miniature microphones that are placed in the user’s ears. The apparatus strives for the closest possible recording of what the user of the tool is hearing while they are taking pictures. This design was chosen in an attempt to enable the later audience to immerse themselves “into the shoes” of the person who originated the content they are experiencing, and to place greater emphasis on the subjective point of view of this original source. (5)

Mali We

chose the African country of Mali as a starting point for thinking about the RAW project because we feel this country has a particularly rich and diverse culture that is not well recognized or understood within Western societies. We conducted a large scale workshop over three weeks in August 2003 in three locations in

Walking Wall (2004) Nell Breyer, Jackbackrack

Cambridge, MA USA

Projection, live video processing 10’ x 4’ on wall and 10’ x 4’ on floorn

A long corridor of black cloth for people to dance along, leads up to a wall of raining feet, eyes and hands. Feet patter along the wall through projections occurring in real and slightly delayed time. Alongside the narrow corridor of black cloth is a white scroll for people to draw on, leave messages and marks on.

Funkenschnorkel (2004) Andrew Brooks

Cambridge, MA USA

Mixed media construction; audio output 15”(w) x 18”(l) x 39”(h) [when not being worn]n

A backpack mounted sound system capable of filling a large space with sound; also LED lighting for illumination and audio peak tracking.

Rabbit Field (2004) Ben Dalton

Cambridge, MA USA

Satin, flannel, PC fans, electronics, Java Approx. 15 x rabbits (Approx. 25 x 35 x 45 cm). A small ad-hoc herd.n

Rabbit Field is an infestation of inflatable rabbit-like forms, filling their display space and inviting tactile interaction. They cover much of the floor, and any other available surfaces, growing in number each night. Each rabbit is self inflating using a simple computer fan, and can sense its internal pressure state by monitoring its fan speed. If a rabbit is squeezed, and partially deflated, the rabbits around it respond, as if out of empathy, deflating themselves. In this way, a wave of deflation ripples out from the squeezed centre.

By connecting an entire field of forms into a network of sensors and output media, interactions between viewer and inflatable are further displayed and amplified as deflation data is passed from one rabbit to the next. The organic feel of the forms and the rhythm of their inflation and deflation in reaction to human touch are easily anthropomorphised by the audience as simple expressions of emotion. This initiates and encourages play and exploration. This piece seeks to encourage and reward a ’tangible dialogue’ between viewer and inflatables, as well as hoping to establish social connection between viewers who co-interact with the system. Rabbit forms were chosen to engage and invite inquiry. These animals also have strong cultural connotations of fertility and innocence, and are prevalent images in modern eastern and western aesthetic. Use of the unique properties of inflatable structures in architecture, art and design has a long and creative history, flirting between chic design and tacky novelty. Funded

in part by a Director’s Grant from the Council for the Arts at MIT.

Cables and Wires (2004) Jeana Frost

Cambridge, MA USA

Mixed method woodcuts Three pieces each 8” x 10”n

Mixed method woodcuts. For these images, original digital photographs are first color separated. These images are then laser cut into wood. The woodblocks are further cut and manipulated by hand and then printed on paper.

Soul o’ Box (2004) Jackbackrack

Cambridge, MA USA

Interactive video sculpture with computer video camera 18”(w) x 18”(l) x 18”(h)n

peering into a box reveals an ever changing intimate small world generated from macro images of self.

Architectural Drip (2004) Brian Knep

Boston, MA USA

Computer, projector, camera 8’(h) x 2’(w)n

An organic pattern grows and flows along one of the gallery pillars. The piece inhabits this structural element and becomes a living architectural detail.

44#444 (2004) Nick Knouf

Cambridge, MA USA

Interactive electronic installation 100cm(h) x 50cm(w) x 50cm(l)n

44#444 is an interactive installation that engages the participant in a reexamination of modern methods of communication. The increased use of cell-phones in the recent past brought about a new form of communication: textmessaging. Most popular overseas, many young people use this as a way to send brief messages to their friends through the phone. What makes text messaging on a phone different from other forms of electronic communication by text is its method of input. Since the majority of cell phones have only 12 keys, and the English alphabet, including both cases numbers, and common symbols, is around 65 characters, a new method of input was devised that took as its first incarnation the use of the letters that already were found on most numeric keypads. For example, in order to enter the letter ”c”, you have to press ”222”; to get ”N”, you have to type ”*66” When you step back and think about this, you can realize how odd of a input method this is. No other common form of input requires you to press the same key multiple times to get a different output. Even capital letters on most keyboards requires you to press two keys at the same time (assuming caps- lock is off). 44#444 makes this disconnect visible by translating short sequences of text into the numeric key presses necessary to enter the text into a common text message. By doing so, 44#444 encourages the participant to consider the unnatural nature of what is by now a rather common mode of communication. The interface to the work consists of an old DEC VT320 terminal. The use of the terminal is an additional way to highlight the disconnect: by using technology from the mid 1980s as an interface to translation of a 21st century mode of communication, the participant is further removed from what is likely a common experience. Participants will be able to take a physical output of their message as printed on thermal paper from 44#444. I will encourage them to write and print out a message that they might send as a text message to a friend and then give the printout to the friend. In this way, I hope to extend the influence of the project outside of the walls of the gallery. Undecided additions to 44#444 might include an audio rendition of the translated message output through headphones. The setup of 44#444 is rather simple. The VT320 terminal and keyboard will sit atop a 50x50cm square box that is about standing height tall. Shortly below the top of the box will be the opening from which the participant can get their printout. I plan to paint the box white. As an example, the string ”The answer to Life, the Universe, and Everything.” is converted to: *84433#166777793377

7#8666#*55544433333 111#84433#*88664448 8833777777733111#16 63#*338883377799984 444466411

Slink (2004) Jeff Lieberman

Cambridge, MA USA

Aluminum, electronics, spring 5’ x 20” x 6”n

A string is made to vibrate at roughly 60 hz by a tubular voice coil [made for the sculpture] - 16 banks of 330 leds each strobe the string at roughly 1spring [which is set to match the resonance of the spring and voice coil] in its fourth mode. Changing the phase of each banks’ leds appears to separate the spring into constituent parts, and changing duty cycle effectively blurs and then erases sections of spring. Various effects are drawn from this initial setup.

Quartet (2003) Andrew Neumann

Boston, MA USA

Wood, LCD screens, cameras, electronics 34” x 17” x 7”n

self referencing wall panel. four cameras monitor the piece as the motors track back and forth across the piece.

Stock Market Skirt (2000) Nancy Paterson

Toronto, Canada

Internet-Interactive Approx. 8’ x 4’n

A blue taffeta and black velvet party dress is displayed on a dressmaker’s mannequin or ’Judy,’ located next to a computer and several monitors of varying sizes. In large type, the stock ticker symbol and price which is being tracked, marches from right to left across the monitor screens as the stock price is continuously updated. Large white numbers and letters on a blue background (matching the blue of the taffeta skirt) scroll in simulation of the pixel board displays used to track stock values on traditional exchange room floor. PERL scripts (running under Linux) extract and analyze stock prices from online stock market quote pages on the internet. These values are sent to a program which determines whether to raise or lower the hemline via a stepper motor and a system of cables, weights and pulleys attached to the underside of the skirt. When the stock price rises, the hemline is raised; when the stock price falls, the hemline is lowered.

Three rotations (2004) James Patten

Cambridge, MA USA

Steel, aluminum, wood, rubber, foam, motors, electronics, computers 3 podiums each measuring about 40”(h) x 24”(w) x 24”(l)

This piece explores three relationships between objects and sound. Three podiums each have a slowly rotating disk, and a scanner which looks at what is on each disk, and translates that into sound. The podiums are placed at the vertices of a triangle on the floor. Walking around this space, the viewer can experience the sound mixing in different ways. The viewer is encouraged to interact with each podium to change the sound it makes. The

first is a surface that is normally grey, but turns black when exposed to water. Visitors moisten a finger with a damp sponge, and then draw on the surface, in effect drawing a melody, which repeats several times, and then fades away. The

second podium has a layer of slow response foam, foam that holds its shape briefly after being squeezed. The visitor squeezes this foam, and the shape of the foam is converted directly into sound. The

third podium has a series of small objects, each of which creates a different sound when it moves under the scanner. Any objects that look different will have different sounds, and gallery visitors are encouraged to explore this space with whatever they may have brought to the exhibition in their pockets.

You’re In Control (Urine Control) (2002) Hayes Raffle, Dan Maynes-Aminzade

Cambridge, MA USA

Urinal, electronics, PIC microcontroller, PC gaming equipment running custom video game 4’ x 8’ x 3’n

The You08re In Control system uses computation to enhance the act of urination. Sensors in the back of a urinal detect the position of a stream of urine, enabling people to play interactive games on a screen mounted above the urinal.

Camera Non-Photo (2004) Ramesh Raskar, Amit Agrawal

Cambridge, MA USA

Multi Flash Camera + Projector + PC 10’ x 10’n

This is a multi-flash camera that can automatically generate stylized images and videos. Strategically positioned flashes cast shadows along silhouettes in the scene. The detected silhouettes are rendered in cartoon style or as technical illustrations. This overcomes the need for per-frame photo editing or 3D scanning of environments.

Specimen (2004) Dan Roe

Cambridge, MA USA

Solar engines, steel Approx. 19” x 19”n

A very unhappy specimen.

Nine Cell Studies (2004) Fran Trainor

Somerville, MA USA

Digital print Each print framed is 10” x 10”, 36” x 36” sq. layoutn

A series of small studies exploring the relationship between cultural and biological production.

Bionic Log (2003) William Tremblay

Allston, MA USA

Wood, steel, latex, electronics Robot is 2’ x 3’ x 6’ and installation is 10’(l) x 6’(w) x 8’(h).n

This is a small-human-sized wood and steel framed pneumatic robot of simple construction and virtually no brains. It also features a video projection of the natural environment the tree parts were found in. It writhes and evokes strong emotional responses from viewers. It is current on display in the Artbots show and is complete, tested, fixable and has not been shown in Boston except at the Allston Open Studios last year.